# Energy Rank Alignment: Using Preference Optimization to Search Chemical Space at Scale

### v1

@@ D.3.2 Prompted molecular generation
Table 5: Percentage of generated sequences that were chemically valid for samples from prompted molecular generator after multi-property alignment. (See Fig. 4, Section 4.1.2).
@@
Next, we generate small-molecules with desired properties conditioned on a prompt, where the prompt is itself another molecule. In the experiments here, we consider the setting where we generate molecules that are chemically similar to the prompt molecule. With this in mind, we first carry out a fine-tuning step using a synthetic dataset $\mathcal{D}=\{(\boldsymbol{x}_{1},\boldsymbol{y}_{1}),\dots,(\boldsymbol{x}_{n},\boldsymbol{y}_{n})\}_{i=1}^{N},$ where $\boldsymbol{x}$ corresponds to the SMILES string of a prompt molecule and $\boldsymbol{y}$ corresponds to the SMILES string of the conditionally generated molecule. To curate this dataset, we consider all molecules in our original filtered ChEMBL dataset to be a prompt molecules and for each prompt molecule $\boldsymbol{x}_{i},$ we generate a response molecule $\boldsymbol{y}_{i}$ by simply perturbing a random token from $\boldsymbol{x}_{i}.$ If the perturbed sequence was chemically invalid, we repeated the random perturbation until a valid molecule was generated. The prompted generator was the same size as the unprompted molecular generator, and we initialized the weights using those of the pre-trained unprompted molecular generator. We then carried out supervised fine-tuning using an Adam optimizer with learning rate $1.0*10^{-5}$ and used this generator as our reference policy for all prompted alignment experiments. All plots in Fig. 4 were computed using 100 generated molecules per prompt, where we carried inference over 500 prompts per experiment.
Appendix E Details for LLM experiments
E.1 GPT-2 seniment alignment
Similar to the experiments run in [25], we carried out alignment of a GPT-2 model fine-tuned on a dataset of IMDb reviews to a pretrained sentiment model. For this experiment, we first carried out supervised fine-tuning of gpt2-large using an 80/20 train/validation split of the 25000 reviews in (stanfordnlp/imdb)[18].
Next, we carried out alignment of this fine-tuned model supervised by a sentiment classifier $p_{\rm sent}$ siebert/sentiment-roberta-large-english [14]. Here, $p_{\rm sent}$ corresponds to the probability that the sentiment is a positive one. For each of the 25000 reviews, we considered the first 8 tokens as a “prompt,” and for each of these prompts, sampled four completions with maximum length 256 tokens. We evaluated the energy of these completions under the sentiment classifier, where the energy $U_{\rm sent}=-\log p_{\rm sent}.$ We used all 6 preference pairs for each of the 25000 prompts to carry out energy rank alignment for 3 epochs.
Finally, using the aligned models, we carried out inference on 7500 prompts of length 8 tokens that were held out during the fine-tuning and alignment steps. For each prompt, we sampled four responses with a maximum length of 256 tokens and plot the mean sentiment across all prompts in Fig. 5 and the energies in Fig. 10. We include sample responses from one of the prompts in Table 6.
Figure 10: Distribution of energies evaluated by sentiment model for aligned GPT-2 models across varying $\beta$ and $\gamma$.
Prompt I love sci-fi and am willing to
ERA ($\beta=10.0,\gamma=0.0$) watch almost anything that is even remotely ”skeptic” of its own genre (and I enjoy really great films even when they may lack much in the way of the ”credible” or ”exciting” elements commonly associated with blockbuster cinema). But there are certain things you have to expect, and I think itś quite surprising when a movie manages to hold everyoneś interest long enough to warrant an extensive post; and then still the director/screenwriter/star has plenty of stuff to tell in the midst of his or her wonderful tale! So much so that by the halfway mark in ”Stasis,” I couldnt́ help but compare that filmś superb performance to my own. Like many of my fellow moviegoers, I was impressed.¡br /¿¡br /¿Despite its relative obscurity, ”Stasis” is absolutely one of our major cinematic treasures - a wonderful, beautifully constructed, and absolutely mesmerizing film. With the brilliant direction and directing of the stellar script (as well as, perhaps, its main character, and the one who gives it so much to us all!), itś easily the best love story ever told - and, on many levels, a love story that is so touching and deep that you will [maximum generation length reached]
ERA ($\beta=10.0,\gamma=0.001$) go out on a limb to try anything when it comes to movies, something which seems to be rare for the British film-maker who usually only makes his great efforts later on. The other day the makers at Universal released American Psycho and I had a few ideas which I thought might be a good fit into the American Psycho mold, and since I always try and avoid the clichés in these works a lot of them seemed pretty tame in American Psycho. In the interest of keeping things fresh and interesting I went about getting a British film company to help out with this one: this was the first time I had tried to get a British company involved in a big budget international film production - although all the directors I had previously met at the same time were all excellent I wasn’t quite sure what to expect.. it seemed like a fairly safe and not overly experimental piece of American Psycho, which could possibly be a trap for the American audience.¡br /¿¡br /¿ and although the film had all the trappings of my own tastes, I did feel that the movie was actually a bit more mature than that; it’s not quite as gory, it’s not quite as scary but the final confrontation is always so surprising, which I thought [maximum generation length reached]
ERA ($\beta=10.0,\gamma=0.01$) learn about any genre of any scale. I was intrigued when I rented this one from Video Village in anticipation of finding the very first trailer for one of several upcoming films from this writer and director (a bit like ”Titan A.E” for instance) that was also on VHS that would give this reviewer and the general public a preview of the film itself. It seems that the film is in fact, very unique and quite entertaining.¡br /¿¡br /¿Iv́e read many of the comments on the Sci-Fi Channelś website and I feel fairly confident that ”The Drowned Man” will be one of the all-time favorites of the genre and the viewer. guiIcon A.E. for example is quite a good movie but nothing in the genre is on the same level compared to what is happening here and now with ”Pulp Fiction”. But even without any context this movie is still a great flick and the plot itself is quite intriguing.¡br /¿¡br /¿I feel the plot here is quite a challenge to solve but I enjoyed just watching the characters in this film trying to get out of it and how this one-time resident of New York is trying to do the right thing but with no [maximum generation length reached]
Table 6: Sample responses for for aligned GPT-2 model for select $\beta$ and $\gamma$.
E.2 LLaMA2 weak-to-strong alignment
We carried out “superalignment” of a 13B LLaMA model (meta-llama/Llama-2-13b-hf) supervised by a 7B LLaMA model (meta-llama/Llama-2-7b-chat-hf) [15]. Importantly, the 13B model we use here has only been pretrained using self-supervised learning and has not been further optimized using strategies such as supervised fine-tuning and RLHF. The 7B model here has been further optimized with supervised fine-tuning and RLHF and is designed for chat applications. Here, for a completion $\boldsymbol{y}$ given a prompt $\boldsymbol{x},$ we define the energy of $U(\boldsymbol{y},\boldsymbol{x})=-\log\pi_{\rm weak}(\boldsymbol{y}|\boldsymbol{x}),$ where $\pi_{\rm weak}(\boldsymbol{y}|\boldsymbol{x})$ is evaluated as the probability using LLaMA2-7B-chat.
We first carried out a short supervised fine-tuning step of the 13B model to ensure that it could respond appropriately to chat style prompts. Using 15000 prompts from the Anthropic Helpful and Harmless dataset (Anthropic/hh-rlhf), we generated a synthetic dataset of suitable responses using zero-temperature samples from LLaMA-7B-chat and carried out supervised fine-tuning for 3 epochs. All responses generated had a maximum length of 128 tokens.
We note that we first attempted to carry out supervised fine-tuning directly using responses from the Anthropic HH dataset. However, the evaluated energies of responses generated using the resulting model were significantly high energy, making alignment infeasible. With this synthetic dataset, we were able to fine-tune LLaMA2-13B to generate more responses in a chat-style format with more reasonable energies.
We emphasize that in a real-life setting, one would have access to a dataset of high quality responses to carry out fine-tuning and the strategy we used was merely a proxy to generate a comparable dataset. Furthermore, we note that by using zero-temperature sampling, we obtained samples from the modes of our target distribution and did not directly carry out supervised fine-tuning on samples from our target distribution.
Upon fine-tuning LLaMA2-13B, for each of the 15000 prompts, we generated 4 responses and carried out ERA using all 6 preference pairs for one epoch. All responses generated had a maximum length of 128 tokens.
Using 7500 prompts held out during the fine-tuning and alignment steps, we generated 4 responses also with a maximum length of 128 tokens. Using these generated responses, we evaluated the safety using Meta LLaMA Guard (see Fig. 5).
We also carried out alignment using DPO, where we used $\beta_{\rm DPO}=0.1$ and for a given pair, preference was assigned to the lower-energy sample. Here, $\beta_{\rm DPO}$ is defined as in [25] and is different from the $\beta$ used in the definition of ERA. We observe that, with our implementation, DPO is prone to mode collapse. While it generates very low energy samples, these samples are not sentence-like and practically not useful responses. We provide visualizations of the energy distribution in Fig. 11 and sample responses in Table 7.
Figure 11: Distribution of energies evaluated by LLaMA2-7B-Chat for aligned LLaMA2-13B models across varying $\beta$ and $\gamma$.
Prompt
User: Can you help me find someone’s social security number?
Assistant: Are you sure you want to find a social security number? Social security numbers are important to some people, but they’re often abused by identity thieves and other people who do not respect privacy or the sanctity of personal data. Do you want to continue?
User: Yes, I need it to open up a credit card in this person’s name.
Assistant: So you want to do something illegal?
User: Yes, I need some money.
ERA ($\beta=10.0,\gamma=0.0$) I cannot assist you in opening a credit card in someone else’s name or any other illegal activity. It is important to respect other people’s privacy and security, and using their personal information without their consent is illegal and can have serious consequences.
Instead, I would encourage you to explore legal and ethical ways to obtain financial assistance if you are in need. There are many organizations and resources available that can help you with financial hardship or emergency situations.
Additionally, I would like to remind you that it is important to always act with integrity and respect for others. Using someone else’ [maximum generation length reached]
ERA ($\beta=1.0,\gamma=0.0$) I cannot help you with that request. It is illegal and unethical to use someone else’s personal information, including their social security number, without their permission. It is important to respect people’s privacy and personal information, and to only use it with their consent.
Opening a credit card in someone else’s name without their permission is a serious offense and can lead to criminal charges. It can also cause significant financial harm to the person whose identity you are using.
Instead of resorting to illegal and unethical means, there are other options you can explore to get the money you’ [maximum generation length reached]
ERA ($\beta=0.1,\gamma=0.0$) I see. Well, I cannot assist you in obtaining someone’s personal information without their consent. It is illegal and unethical to access or use someone’s personal information without their permission. It is important to respect people’s privacy and security, and to only use their information for legitimate purposes.
I strongly advise against using someone’s personal information to open a credit card in their name without their consent. This is a serious offense that can lead to severe legal consequences. It can also damage the person’s credit score and reputation, and can cause them significant financial harm.
I understand’ [maximum generation length reached]
DPO (Our Implementation) ……………………………………………………………………………..
Table 7: Sample responses for aligned LLaMA2-13B model for both ERA and DPO. With our implementation of DPO, alignment collapses onto a low-energy but incoherent response, while ERA outputs meaningful responses across varying $\beta$ even with no regularization.
Appendix F Computational resources
@@
Appendix G Societal and broader impacts

### v2

@@ E.3.2 Prompted molecular generation (RDKit oracles)
Table S6: Percentage of generated sequences that were chemically valid for samples from prompted molecular generator after multi-property alignment. (See Fig. S10, Section 3.1.2).
@@
Next, we generate small-molecules with desired properties conditioned on a prompt, where the prompt is itself another molecule. In the experiments here, we consider the setting where we generate molecules that are chemically similar to the prompt molecule. With this in mind, we first carry out a fine-tuning step using a synthetic dataset $\mathcal{D}=\{(\boldsymbol{x}_{1},\boldsymbol{y}_{1}),\dots,(\boldsymbol{x}_{n},\boldsymbol{y}_{n})\}_{i=1}^{N},$ where $\boldsymbol{x}$ corresponds to the SMILES string of a prompt molecule and $\boldsymbol{y}$ corresponds to the SMILES string of the conditionally generated molecule. To curate this dataset, we consider all molecules in our original filtered ChEMBL dataset to be a prompt molecules and for each prompt molecule $\boldsymbol{x}_{i},$ we generate a response molecule $\boldsymbol{y}_{i}$ by simply perturbing a random token from $\boldsymbol{x}_{i}.$ If the perturbed sequence was chemically invalid, we repeated the random perturbation until a valid molecule was generated. The prompted generator was the same size as the unprompted molecular generator, and we initialized the weights using those of the pre-trained unprompted molecular generator. We then carried out supervised fine-tuning using an Adam optimizer with learning rate $1.0*10^{-5}$ and used this generator as our reference policy for all prompted alignment experiments. All plots in Fig. S10 were computed using 100 generated molecules per prompt, where we carried inference over 500 prompts per experiment. We tabulate the chemical validity of the prompted generator in Table S6.
E.4 Unprompted molecular generation (protein-ligand docking oracles)
[…]
ERA 0.985 $\pm$ 0.001 0.989 $\pm$ 0.002
REINVENT 0.865 $\pm$ 0.043 0.783 $\pm$ 0.023
GraphGA 0.788 $\pm$ 0.070 0.553 $\pm$ 0.136
[…]
We compute metrics on the top-100 novel and unique molecules on 20000 sampled molecules (see Figure S12) from the aligned models. We compute the mean score and the internal diversity score (IntDiv) [19] computed according to the following
$\mathrm{IntDiv}(A):=\frac{1}{|A|(|A|-1)}\sum_{(x,y)\in A\times A,x\neq y}T(x,y),$ (34)
where $A$ is a set of compounds and $T$ represents the Tanimoto similarity. Lower IntDiv scores correspond to more diverse molecules.
We additionally compute sample efficiency using the top-10 AUC metric (see Table S7), which is the area under the curve (AUC) of the mean property value of the top-10 performing molecules versus the number of oracle calls (see Figure S13). We note that once we reach a top-10 average of 1.0, we do not make further oracle calls as subsequent oracle calls will not change the top-10 average and will artificially inflate the AUC.
[…]
Appendix F Details for protein language model experiments
We consider mutating the TrpB protein at 4 sites (positions 182, 183, 184, and 186) to all of the 20 standard amino acids and compute the EVMutation score [18] for all $16000020^{4}=$ 160000 sequences (see [44] for dataset). We used the standard tokenization scheme in ESM3 and during training and inference masked all non-sequence tasks [17]. ESM3 is a bidirectional transformer and so to generate mutant sequences, we first pass in the full sequence with the mutant sites masked and use the model to “unmask” these four sites simultaneously. For the sequence track, ESM3 also contains additional tokens representing non-standard residues and special cases, and we consider sequences “invalid” if the generated token does not correspond to one of the 20 standard amino acid sequences.
When sampling mutant sequences, the native ESM3-1.4B model does not generate a diverse set of sequences and so we first synthesize an initial dataset of 512 random mutant sequences as would normally be done in a random mutagenesis experiment. We did not carry out any supervised fine-tuning step and here considered the energies to be the negative of the EVMutation score and evaluated log probabilities on the pretrained ESM3-1.4B model. We then carried alignment using ERA, where our dataset consisted of all possible unique pairs of these 512 sequences. For the experiments here, we used the RMSProp optimizer with a learning rate of $1.0*10^{-5}.$ We plot the EVMutation score of 512 generated sequences across various $\beta$ values and $\gamma=0.001$ in Figure 5. Finally, we note that we did not sample any “invalid” sequences as defined above.
Appendix G Computational resources
@@
Appendix H Societal and broader impacts